# Constant-Size Certificates for Leader Election in Chordal Graphs and Related Classes

### v1

@@ 4.2 Spanning Tree
We present a certificate verifying the existence of a spanning tree in any dismantlable labeled graph $G=(V,E)$. The labeling function assigns each $v\in V$ a label in $L=\{0,1\}$, with exactly one node labeled 1 (the root $r$) and all others labeled 0. We define $C=\{0,1,2\}$ to be the certificate and each node is assigned a value from $C$ by $x_{C}:V\rightarrow C$. The following must hold for every node $v$ of $G$ with value $i$.
@@
    Node $v\neq r$ has at least one neighbor with value $(i-1)\mod 3$.
@@
We show that if N1–N2 are satisfied for every node in $G$, adding an edge from any node $v$ distinct from the root with value $i$ to any neighbor of $v$ with value $i-1\mod 3$, induces a spanning tree.
@@
In any graph $G$ there exists a certificate assignment that satisfies N1–N2 for every node in $G$.
@@
Let $T$ be a spanning subgraph of $\vec{G}$ in which each node $v$, except the root, arbitrarily selects a parent among its neighbors with value equal to $(\textit{value of }v-1)\mod 3$. We show that $T$ is a tree. Since $\vec{G}$ is acyclic by Lemma 4.7, it follows that $T$ is also acyclic. We prove that $T$ is connected.
@@
Let $u$ be a node in $T$. Since from Lemma 4.7, $\vec{G}$ and as a result $T$ does not contain cycles, the directed path starting at $u$ must end at a node without a successor in $T$. From N1, the root is the only node without a smaller neighbor. Hence, any node $u$ must be in a path that reaches the root, and as a result $T$ is connected.
Finally, since every node in $G$ has a label and must be in $T$, from Lemmas 4.7 and 4.9 we obtain the following:
@@
5 From Certificate to Self–Stabilization
An algorithm is called self-stabilizing when it can autonomously recover from transient faults. In particular, a distributed self-stabilizing algorithm is usually defined as a set of rules evaluated at every node. Each rule is of the form:
@@ 5 From Certificate to Self–Stabilization
where guard is a boolean predicate over the variables of the closed neighborhood of the node and the command specifies how the node updates its variables. Here, we assume that each node only updates its state variable. A self-stabilizing algorithm is said to be silent if after some finite time no node updates its variables anymore. When activated, each node can perform computations based on its state and the states of its neighbors and update its own state. The order of node activations is determined by a scheduler. A scheduler is characterized both by synchronicity and fairness. In particular, here we consider an asynchronous scheduler meaning that at each step, an arbitrary (non-empty) subset of nodes may be activated. In terms of fairness, we assume a Gouda fair scheduler [Gouda2001theory]. According to [dubois2011taxonomy], Gouda fairness is the strongest fairness assumption. It requires that for any configuration that appears infinitely often during the execution of an algorithm, every configuration that is reachable by executing any non-empty subset of enabled commands (i.e., commands for which the guard is true) also appears infinitely often. We assume that the graphs we consider are finite. In this setting, Gouda fairness ensures that any configuration that is infinitely often reachable is eventually reached. In this section we show how to obtain a silent self–stabilizing algorithm that works under a Gouda fair scheduler, given a certification scheme.
We say that under a given algorithm $A$, a certificate assignment is final if no node satisfies any of the guards in $A$. We define a correct final certificate assignment to be one that is both final with respect to an algorithm $A$ and correct with respect to a certification scheme. Call the set of states in a given certificate $C$. Informally, every node that detects that a condition of the certification scheme is violated sets its state to a reset state that we denote $\bot$. Additionally, every node that has at least one neighbor that has detected an error and reset its state, also changes its state to the reset state. Hence, an error detection is propagated throughout the graph from node to neighboring node. When a node that is at the reset state is activated, it non-deterministically switches to a state from $C$. Gouda fairness ensures that the graph eventually reaches a correct global configuration.
A node executing Algorithm 1 evaluates the guards in order. If the first guard is enabled, the node performs the corresponding command and deactivates without evaluating the second guard.
Input: The certification scheme $(C,R)$ and a certificate assignment $x_{C}:V\rightarrow C$
@@
/* states from $C$ are always chosen non-deterministically */
@@
S2 : v.state = $\bot$ $\rightarrow$ v.state $\leftarrow$ state from $C$
Algorithm 1 General Certificate to Self-Stabilizing Algorithm
@@
Let $(C,R)$ be a certification scheme satisfying the Existence and Soundness conditins for some task $\mathcal{T}$ in a connected graph $G$. Starting from any arbitrary configuration, any Gouda fair execution of Algorithm 1 on $G$ stabilizes to a correct final configuration.
@@
Let $G^{*}$ be a configuration where no node satisfies the guard of S1 or of S2, that is, $G^{*}$ is a final correct configuration. By definition, $\mathcal{T}$ is solved in $G^{*}$, since $(C,R)$ is a certification scheme for $\mathcal{T}$. Suppose for contradiction that there exists a configuration $G^{\prime}\neq G^{*}$ that is infinitely often reachable during the execution of Algorithm 1. Since $G^{*}$ is final, it cannot be reachable from $G^{\prime}$. As $G^{\prime}$ is not final, either $G^{\prime}=G_{\bot}$ (i.e., all nodes are in state $\bot$) or the guard of S1 is enabled for some node. In the former case, $G_{\bot}$ trivially appears infinitely often. In the latter, by S1, there exists an activation sequence that leads to $G_{\bot}$. By Gouda fairness, for every a configuration that appears infinitely often, every reachable configuration also appears infinitely often. Hence, in either case $G_{\bot}$ appears infinitely often. Due to S2, once all nodes in $G_{\bot}$ are activated, the configuration becomes $G^{*}$, contradiction. Thus, no configuration $G^{\prime}\neq G^{*}$ can appear infinitely often.
A self–stabilizing algorithm can be obtained by combining the respective certification scheme of each of the previous sections with Algorithm 1.
@@ 6 Discussion and Open Problems
In Sections˜3 and 4 we gave certification schemes for leader election, with constant size per edge. A natural direction for further research is studying whether our certification scheme from Section˜4 also works for dismantlable graphs. In particular, only the proof of ˜2.2 depends on the graph being $K_{4}$-free. However, removing this restriction has been very challenging. If this is not possible, another direction is considering bridged or Helly graphs, which are dismantlable but enjoy interesting properties that could be useful for addressing our problem. Alternatively, if our certification scheme can be proven to also work on dismantlable graphs, it is worth exploring other relevant classes. For example, [fieux2020hierarchy] gives a hierarchy on classes (based also on the work of [mazurkiewicz2007locally]) that have dismantling orderings, called s-dismantlable, where 1-dismantlable graphs correspond to the dismantlable graphs defined in Section 4. Informally, for $s>1$, [fieux2020hierarchy] defines a node to be $s+1$ dismantlable if the graph induced by its open neighborhood is $s$-dismantlable and a graph to be $s$-dismantlable if it is reducible to a single vertex by successive deletions of $s$-dismantlable vertices. These directions for future research are summarized in Figure˜4.
@@
Figure 4: A graph class hierarchy, where each arrow represents an inclusion relation. The classes highlighted in green are an easy extension of [chalopin2024stabilising] and those in yellow are considered in this paper for leader election certification with constant memory per edge.
@@
Another potential direction for future research is exploring whether it is possible to automatically transform a certificate into a self-stabilizing algorithm with a weaker scheduler, as defined in the taxonomy of [dubois2011taxonomy], even using non-constant memory. For example, the transformation from [BlinFP14], which uses logarithmic memory but works even with an unfair scheduler, contrasts with our algorithm, which uses constant memory but requires a Gouda fair scheduler. These two results suggest that there may be a trade-off between the strength of the scheduler and the memory requirements that has not yet been fully explored.

### v2

@@ 4.2 Spanning Tree
We present a labeling constructing a spanning tree in any dismantlable graph $G=(V,E)$, with inputs. The input function assigns each $v\in V$ an input in $I=\{0,1\}$, with exactly one node receiving 1 (the root $r$) and all others receiving 0. We define $L=\{0,1,2\}$ to be the set of labels and each node is assigned a value from $L$ by $x_{L}:V\rightarrow L$. The following must hold for every node $v$ of $G$ with label $l$.
@@
    Node $v\neq r$ has at least one neighbor with label $(l-1)\mod 3$.
@@
We show that if N1–N2 are satisfied for every node in $G$, adding an edge from any node $v$ distinct from the root with label $l$ to any neighbor of $v$ with label $l-1\mod 3$, induces a spanning tree.
We first prove the Existence property.
@@
In any graph $G$ there exists a labeling that satisfies N1–N2 for every node in $G$.
@@
Let $T$ be a spanning subgraph of $\vec{G}$ in which each node $v$, except the root, arbitrarily selects a parent among its neighbors with label equal to $(\textit{label of }v-1)\mod 3$. We show that $T$ is a tree. Since $\vec{G}$ is acyclic by Lemma 4.7, $T$ is also acyclic. We prove $T$ is connected.
@@
Let $u$ be a node in $T$. Since from Lemma 4.7, $\vec{G}$ and as a result $T$ does not contain cycles, the directed path starting at $u$ must end at a node without a successor in $T$. From N1, the root is the only node without a successor. Hence, any node $u$ must be in a path that reaches the root, and as a result $T$ is connected.
Since every node in $G$ has a label and must be in $T$, from Lemmas 4.7 and 4.9 we obtain:
@@
5 From Labeling to Self–Stabilization
A distributed self-stabilizing algorithm can autonomously recover from transient fault and is usually defined as a set of rules evaluated at every node. Each rule is of the form:
@@ 5 From Labeling to Self–Stabilization
where guard is a boolean predicate over the variables of the closed neighborhood of the node and the command specifies how the node updates its variables. Here, we assume that each node only updates its state variable. A self-stabilizing algorithm is said to be silent if after some finite time no node updates its variables anymore. When activated, each node can perform computations based on its state and the states of its neighbors and update its own state. The order of node activations is determined by a scheduler. A scheduler is characterized both by synchronicity and fairness. In particular, here we consider an asynchronous scheduler meaning that at each step, an arbitrary (non-empty) subset of nodes may be activated. In terms of fairness, we assume a Gouda fair scheduler [Gouda2001theory]. According to [dubois2011taxonomy], Gouda fairness is the strongest fairness assumption. It requires that for any configuration that appears infinitely often during the execution of an algorithm, every configuration that is reachable by executing any non-empty subset of enabled commands (i.e., commands for which the guard is true) also appears infinitely often. We assume that the graphs we consider are finite. In this setting, Gouda fairness ensures that any configuration that is infinitely often reachable is eventually reached. In this section we show how to obtain a silent self–stabilizing algorithm that works under a Gouda fair scheduler, given a labeling scheme.
We say that under a given algorithm $A$, a label assignment is final if no node satisfies any of the guards in $A$. We define a correct final label assignment to be one that is both final with respect to an algorithm $A$ and correct with respect to a labeling scheme. Call $L$ the set of states in a given labeling. Informally, every node that detects a condition of the labeling scheme being violated sets its state to a reset state that we denote $\bot$. Every node that has at least one neighbor in a reset state, also changes its state to the reset state. Hence, an error detection is propagated throughout the graph from node to neighboring node. When a node that is at the reset state is activated, it non-deterministically switches to a state from $L$. Gouda fairness ensures that the graph eventually reaches a correct global configuration.
A node executing Algorithm 1 evaluates the guards in order. If the first guard is enabled, the corresponding command is performed and the second guard is not evaluated.
Given a labeling scheme $(L,R)$, a graph $G$ and a vertex $v\in G$, a possible state for $v$ in $L$ is a state $s$ from $L$ such that there exists a 1-ball $(B,x)\in R$ where the center of $B$ is labeled by $x(v)=s$ and $B$ is isomorphic to $B_{1}(v)$. If $(L,R)$ is a certification scheme of finite type, then the set of possible states for each node in any graph $G$ is finite.
Input: The labeling scheme $(L,R)$ and a label assignment $x_{L}:V\rightarrow L$
@@
/* states from $L$ are always chosen non-deterministically */
@@
S2 : v.state = $\bot$ $\rightarrow$ v.state $\leftarrow$ possible state from $L$
Algorithm 1 General Labeling to Self-Stabilizing Algorithm
@@
Let $(L,R)$ be a finite labeling scheme of finite type satisfying the Existence and Soundness conditions for some task $\mathcal{T}$ in a connected graph $G$. Starting from any arbitrary configuration, any Gouda fair execution of Algorithm 1 on $G$ stabilizes to a correct final configuration.
@@
Let $G^{*}$ be a configuration where no node satisfies the guard of S1 or of S2, that is, $G^{*}$ is a final correct configuration. By the soundness property of $(L,R)$ for $\mathcal{T}$, $\mathcal{T}$ is solved in $G^{*}$. Consequently, in any final configuration, $\mathcal{T}$ is solved.
We now show that any execution eventually reaches a final configuration. Since $(L,R)$ is of finite type, for each node $v$, the set of possible states is finite and thus in any execution, every node $v$ enters only in a finite number of different states. Since $G$ is finite, the number of different configurations existing in any execution (finite or infinite) is also finite. By the existence property of $(L,R)$ for $\mathcal{T}$ in $G$, there exists a configuration $G^{*}=(G,x_{L})$ of $G$ such that in $G$, every 1-ball $(B_{1}(v),x_{L}|_{N[v]})$ belongs to $R$. Suppose there exists an infinite execution of the algorithm. This execution does not reach any final configuration, i.e., at each step, there is a node that can execute S1 or S2. In the second case, it means that there is node that is labeled by $\bot$. In the first case, it means that one can reach in one step a configuration containing at least one node labeled $\bot$. Starting from a configuration containing one node labeled $\bot$, there exists a sequence of activation of nodes that enables reaching the configuration $G_{\bot}$ where all nodes are in state $\bot$. Then by Gouda fairness, in any infinite execution, the configuration $G_{\bot}$ is reached infinitely often. From $G_{\bot}$, there exists a sequence of node activations that enables reaching the configuration $G^{*}$. Again, by Gouda fairness, this implies that any infinite execution eventually reaches $G^{*}$. Since $G^{*}$ is a final configuration where no node can be activated, we obtain a contradiction with the existence of an infinite execution.
A self–stabilizing algorithm can be obtained by combining the respective labeling scheme of each of the previous sections with Algorithm 1.
@@ 6 Discussion and Open Problems
In Section˜4 we gave a labeling scheme that implies a unique sink defined to be a leader, with labels encoding the orientation of all edges incident to every node. Observe that with our labeling schemes for leader election in chordal graphs and in trees, any node in the graph can be the sink and thus designated as a leader. Indeed, any chordal graph $G$ contains at least two simplicial vertices, and thus for any vertex $v$, there exists a simplicial ordering that ends at $v$. By Section˜3.2, there exists then an orientation of $\vec{G}$ satisfying C1–C3 where $v$ is the unique sink of $\vec{G}$. This property is however no longer true for dismantlable graphs. Consider for example the graph $G$ of Figure 3 and observe that $a$ is the only dominated vertex of $G$. This implies that $a$ cannot be the sink of any orientation satisfying D1–D3 obtained by Section˜4.1. In fact, one can show that in any orientation $\vec{G}$ of $G$ satisfying D1–D3, $a$ is not a sink. Indeed, suppose there exists such an orientation. From D2 at $b$ and $c$, the edges $bd$ and $ce$ should be oriented as $\overrightarrow{db}$ and $\overrightarrow{ec}$. Consider now the edge $de$ and assume without loss of generality that it is oriented as $\overrightarrow{de}$. Then $d$ has two non-adjacent outgoing neighbors $b$ and $e$ that do not have a common neighbor, and thus D2 does not hold at $d$. We believe however that our labeling scheme can be easily modified to elect any node in the graph, by adding an extra component to the labels of the nodes in order to encode a directed path from the unique sink to the designated leader.
A natural direction for further research is studying whether our labeling scheme from Section˜4 ensures the existence of a unique sink in any dismantlable graph. In particular, only the proof of ˜2.2 depends on the graph being $K_{4}$-free. However, removing this restriction has been very challenging. One can also try to adapt to ($K_{4}$-free) dismantlable graphs the approach used in [ChChKo-meshed] to design labeling schemes for leader election in meshed graphs. However, this also seems very challenging, as it relies heavily on properties of the distance function in meshed graphs, and dismantlable graphs do not enjoy similar properties (even in the $K_{4}$-free case).
@@
If one can find a labeling scheme for leader election in dismantlable graphs, it is worth exploring other relevant classes. For example, [fieux2020hierarchy] gives a hierarchy on classes (based also on the work of [mazurkiewicz2007locally]) that have dismantling orderings, called s-dismantlable, where 1-dismantlable graphs correspond to the dismantlable graphs. Informally, for $s>1$, [fieux2020hierarchy] defines a node to be $s+1$ dismantlable if the graph induced by its open neighborhood is $s$-dismantlable and a graph to be $s$-dismantlable if it is reducible to a single vertex by successive deletions of $s$-dismantlable vertices. For a summary, see Figure˜4.
@@
Figure 4: A graph class hierarchy, where each arrow represents an inclusion relation. The green classes are the result of [chalopin2024stabilising] and subclasses of the graphs in this paper that generalize the result of [chalopin2024stabilising]. Yellow classes are considered in this paper. Blue classes are considered in [ChChKo-meshed].
@@
Finally, a potential direction for future research is exploring whether it is possible to automatically transform a labeling into a self-stabilizing algorithm with a weaker scheduler, as defined in the taxonomy of [dubois2011taxonomy], even using non-constant memory. For example, the transformation from [BlinFP14], which uses logarithmic memory but works even with an unfair scheduler, contrasts with our algorithm, which uses constant memory but requires a Gouda fair scheduler. These two results suggest that there may be a trade-off between the strength of the scheduler and the memory requirements that has not yet been fully explored.